OpenStudy (anonymous):

can some one help me put please????? solve for x: -3x + 2 = 23

OpenStudy (bahrom7893):

-3x = 23 - 2 -3x = 21 x = -7

OpenStudy (anonymous):

can you help me out with this one to please? its is solve for x : 9x + 4 = 8x + 5

OpenStudy (bahrom7893):

x=1

OpenStudy (anonymous):

thank you for helping me

OpenStudy (bahrom7893):

np

OpenStudy (anonymous):

wat about this one ? it the same type of problem 4( x + 3 ) = 12

OpenStudy (bahrom7893):

same thing x = 0

OpenStudy (anonymous):

ok I only need help with these last two 8 ( x + 3 ) = 7 ( x + 3) and 3 ( x - 2) + 5 ( x + 3) = 33